OpenStudy (anonymous):

So....? Switching to which base would allow you to easily reduce this down to the fraction 3/2 ? log_4(8) a) base 8 b) base 4 c) base 2 d) base 10

OpenStudy (campbell_st):

well base 2... since 4 = 2^2 and 8 = 2^3 the change of base formula is \[\log_{a}(b) = \frac{\log(b)}{\log(a)}\]